Bawa Selesa Bhd, has enjoyed substantial economic profit derived from patents covering the manufacturing of a special shock absorber. Market demand and marginal revenue relations for the shock absorber are:
P=5000–0.05Q
MR = MTR/MQ = 5000 -0.1Q
Fixed costs are nil, because research and development expenses have been fully amortized during previous periods. Average variable costs are constant at RM4, 000 per unit
a) Calculate the profit-maximizing price/output combination and economic profits if Bawa Selesa enjoys an effective monopoly on the shock absorber market due to its patent protection
1
Expert's answer
2017-04-05T14:31:07-0400
P=5000–0.05Q,
MR = MTR/MQ = 5,000 - 0.1Q
FC = 0, AVC = RM4,000 per unit.
a) The profit-maximizing price and output combination is in the point where MR = MC = P, so:
MC = TC' = (FC + VC)' = (0 + 4,000Q)' = 4,000.
5,000 - 0.1Q = 4,000,
Q = 10,000 units.
P = MR = 5,000 - 0.1*10,000 = RM4,000.
Economic profits are:
TP = (P - ATC)*Q = (4,000 - 4,000)*10,000 = 0.